OpenStudy (watermelon14):

find the intersection of the following graphs over the interval [0,2pi). f(x)=sin2x g(x)=cosx

OpenStudy (watermelon14):

@jdoe0001 do i set them equal to each other?

OpenStudy (freckles):

solve f(x)=g(x) for x that is solve sin(2x)=cos(x) use double angle identity for sin(2x)

OpenStudy (jdoe0001):

yeap, you set f(x) = g(x) whatever value "x" gets, is where they intersect

OpenStudy (jdoe0001):

be wary of cancelling cosine in the simplification, btw

OpenStudy (watermelon14):

ok so does it look like 2sinxcosx=cox

OpenStudy (watermelon14):

how do i simplify that down further?

OpenStudy (jdoe0001):

so.... \(\begin{array}{llll} f(x)=sin(2x)\\ g(x)=cos(x) \end{array}\implies sin(2x)=cos(x) \\ \quad \\ {\color{brown}{ 2sin(x)cos(x) }}=cos(x)\implies 2sin(x)cos(x)-cos(x)=0 \\ \quad \\ cos(x)[2sin(x)-1]=0\implies \begin{cases} cos(x)=0\implies &x=cos^{-1}(0)\\ 2sin(x)-1=0\implies &x=sin^{-1}\left( \frac{1}{2} \right) \end{cases}\)

OpenStudy (watermelon14):

ohhhh thanks!

OpenStudy (jdoe0001):

you could have cancelled cosine in the simplification, instead of doing common factor but you'd have lost a valid value in the process tis what I meant, be wary of cancelling it :)